OpenStudy (anonymous):

Assume it takes 7.00 minutes to fill a 30.0-gal gasoline tank. A. Calculate the rate at which the tank is filled in gallons per second. B. Calculate the flow rate of the gasoline in cubic meters per second. C. Determine the time interval, in hours, required to fill a 1.00m^3 volume at the same rate. (1US gal= 231 in.^3)

OpenStudy (amistre64):

\[\frac{g}{m}\frac{m}{s}=\frac{g}{s}\]

OpenStudy (amistre64):

\[\frac gs \frac {m^3}g=\frac {m^3}s\]

OpenStudy (anonymous):

what is the solution ?

OpenStudy (amistre64):

youd have to refer to conversion equivalents to determine that ... but once you know the conversions .. the math is simple enough

OpenStudy (amistre64):

how many seconds in a minute? how many cubic meters in a gallon? etc .....

OpenStudy (amistre64):

looks like for gallons to meters^3 you would need to determine the conversion of inches to meters ...

OpenStudy (anonymous):

what is m?

OpenStudy (anonymous):

can you solve g/m m/s = g/s